OpenStudy (anonymous):

I need help write hyperbola equation 4x^2-26y^2+10=0

OpenStudy (anonymous):

\[(x ^{2}/a ^{2})-(y ^{2}/a ^{2)}=1\]

OpenStudy (anonymous):

\[4x^2-26y^2=-10\]

OpenStudy (anonymous):

\[\frac{4x^2-26y^2=-10}{-10}\]

OpenStudy (anonymous):

The should be no coefficient (or, rather, a 1 in front of x^2 and y^2. You can leave the 10 on the LHS for the moment, so you don't have to deal with negative. But first divide by (4)(26)

OpenStudy (anonymous):

yes,but is decimal right?

OpenStudy (anonymous):

frist division by 4?

OpenStudy (anonymous):

\[x^2-\frac{26y^2}{4}=\frac{-10}{4}\]?

OpenStudy (anonymous):

I use too no problem to write equation, but this number seem not even, Make me confuse

OpenStudy (anonymous):

26/4 is decimal

OpenStudy (anonymous):

move 10 to left hand side,than (4*26)+(-10) ?

OpenStudy (anonymous):

Yes, the numbers are not coming out like squares, it usually works out. Make sure you wrote down the problem right.

OpenStudy (anonymous):

yes, I did

OpenStudy (anonymous):

I never see like dis problem before,I will send you the page,check number 9

OpenStudy (anonymous):

Well, the question doesn't say put in form of hyperbola. It asks is it parabola, circle, ellipse, or hyperbola.

OpenStudy (anonymous):

but I write befor graph

OpenStudy (anonymous):

I can state is hyperbola, but I need to know x and y I can graph

OpenStudy (anonymous):

Yeah, it is a hyperbola, but the numbers are complicated. I graphed it on wolfram http://www.wolframalpha.com/input/?i=4x^2-26y^2%2B10%3D0

OpenStudy (anonymous):

ty